OpenStudy (anonymous):

Match each expression to its approximate value to the nearest tenth. 3+√5 √7+2 13-√1o a)9.8 b)4.6 c)5.2

OpenStudy (anonymous):

\[3+\sqrt{5} ---- c\]\[\sqrt{7}+2 -----b\]\[13-\sqrt{10}------a\]